OpenStudy (mervatm):

please help will medal and fan. its multiple choice The approach ramp used by a daredevil motorcyclist for flying over a collection of flaming tires has a rise of 44 feet for every 80 feet in horizontal distance. Find the grade of the ramp. Round to the nearest whole percent. 0.55% 60% 18% 55%

OpenStudy (anonymous):

Grade is Rise/Run

563blackghost (563blackghost):

Grade: \(\Huge{Grade=\frac{rise}{run} \times 100}\) The rise is \(\large{44}\) feet and the run is \(\large{80}\) feet..so we input this... \(\Huge{Grade=\frac{\color{red}{44}}{\color{red}{80}} \times 100}\) Simplify... \(\Huge{{\color{red}{.55} \times 100}}\) So multiply this and you'll get your answer ^^